### v1

@@ 3.2 Phase-averaged Spectral analysis
When comparing these results with Díaz Trigo’s estimates of at 5 kpc and at 10 kpc in the 0.6-10 keV range, we observe that our findings are in line with these estimates, particularly when considering the range of distances. Specifically, our luminosity results suggest that the pulsar is relatively bright at the higher distance of 9.2 kpc, which aligns with the higher end of the luminosities reported by Díaz Trigo. Additionally, the decreasing luminosities associated with shorter distances emphasize the inverse relationship between distance and brightness.
@@
These results also underscore the critical role of accurate distance measurements in astrophysical studies, as variations in distance can significantly affect luminosity calculations. The discrepancies between our luminosity estimates and those reported by Boirin et al. at different distances may stem from differences in energy ranges analyzed or variations in the flux values used in their calculations. Furthermore, our findings highlight the importance of using multiple distance estimates to obtain a comprehensive understanding of the source’s intrinsic properties, as the luminosity values can provide insights into the physical processes occurring in EXO 0748-676, particularly during burst events. Overall, the luminosity derived from our analysis demonstrates that EXO 0748-676 remains a key target for understanding the behavior of low-mass X-ray binaries, contributing to the ongoing discourse regarding the physical mechanisms at play during X-ray outbursts.
@@ 4 Conclusion
In conclusion, the timing analysis of X-ray bursts from the neutron star system provides valuable insights into the interplay of thermonuclear processes, accretion dynamics, and magnetic field interactions. The primary burst at signifies a substantial thermonuclear explosion on the neutron star’s surface, marked by a sharp increase in count rate, indicating a rapid energy release. Following this major burst, smaller events, such as the recoil burst at , illustrate the system’s stabilization after the initial explosion. The alignment of burst timings with the neutron star’s optical period suggests a possible modulation by the star’s rotation and periodic accretion, revealing how these factors influence the timing and intensity of bursts. Additionally, the analysis across different energy bands highlights the relationship between energy levels and burst characteristics, with soft X-rays linked to cooling phases, mid-energy emissions representing peak explosions, and hard X-rays associated with thermonuclear ignition. Overall, these findings reinforce our understanding of the mechanisms behind Type I X-ray bursts and pave the way for future research into the physics governing accretion and burst phenomena in neutron star systems, enhancing our knowledge of stellar evolution and the extreme conditions in the universe.
The spectral analysis of EXO 0748-676 conducted using the NuSTAR observatory’s FPMA and FPMB instruments in the 3-79 keV energy range provides significant insights into the source’s emission characteristics, revealing a photon index of and a cutoff energy of , which suggests the presence of a hot corona around the neutron star and indicates the nature of the emission processes at play. The measured flux of approximately highlights the variable nature of accretion-driven systems. Comparative analysis with earlier observations from the IBIS/ISGRI instrument on the INTEGRAL satellite further enhances our understanding of the source’s behavior over time, as the slightly softer photon indices and higher cutoff energies observed in 2003 and 2004 suggest that the emission may have been influenced by changes in coronal temperature and accretion dynamics during those periods. The variability in flux measurements across the different observational epochs underscores the dynamic nature of the source, reflecting typical fluctuations seen in accreting neutron stars. Overall, this analysis illustrates the effectiveness of high-energy spectral fitting in probing the physical processes governing the emission from EXO 0748-676, contributing to the broader understanding of neutron star physics and the complexities of accretion in binary systems, while the consistency of spectral parameters with previous studies also validates the reliability of the methods employed, establishing a foundation for future investigations into similar astrophysical objects.
Additionally, our analysis of the luminosity of EXO 0748-676, derived from various distance estimates and flux measurements, provides critical insights into the characteristics of this low-mass X-ray binary. The calculated luminosities vary based on distance, highlighting the importance of accurate distance measurements in astrophysical studies. Notably, our results align well with previous estimates from Díaz Trigo and Boirin et al., suggesting that EXO 0748-676 exhibits significant intrinsic brightness, particularly at the higher distance of 9.2 kpc. These findings emphasize the need for continued investigation into the physical processes governing X-ray bursts in this system, as understanding the luminosity and behavior of such sources is essential for elucidating the mechanisms driving X-ray emission in low-mass X-ray binaries. Ultimately, this study contributes to a more comprehensive understanding of EXO 0748-676 and its role in the broader context of X-ray astronomy.

### v2

@@ 3.2 Phase-averaged Spectral analysis
When comparing these results with Díaz Trigo’s estimates of at 5 kpc and at 10 kpc in the 0.6-10 keV range, we observe that our findings are in line with these estimates, particularly when considering the range of distances. Specifically, our luminosity results suggest that the pulsar is relatively bright at the higher distance of 9.2 kpc, which aligns with the higher end of the luminosities reported by Díaz Trigo.
@@
The luminosities derived from both NuSTAR models are consistent with previous studies when taking into account the different distance estimates and the variability in the accretion environment of EXO 0748-676. The slightly higher luminosity values obtained from Model 1 reflect the marginally higher fluxes observed with this model, while Model 2 introduces greater uncertainty in the results due to its broader error margins. Overall, the results provide valuable insight into the emission properties of EXO 0748-676 and reinforce the importance of accurate distance measurements in determining the physical characteristics of such systems, as variations in distance can significantly affect luminosity calculations. The discrepancies between our estimates and those reported by Boirin et al. at different distances may stem from differences in energy ranges analyzed or variations in the flux values used in their calculations. Furthermore, our findings highlight the importance of using multiple distance estimates to obtain a comprehensive understanding of the source’s intrinsic properties, as the luminosity values can provide insights into the physical processes occurring in EXO 0748-676, particularly during burst events. Overall, the luminosity derived from our analysis demonstrates that EXO 0748-676 remains a key target for understanding the behavior of low-mass X-ray binaries, contributing to the ongoing discourse regarding the physical mechanisms at play during X-ray outbursts.
@@ 4 Conclusion
The analysis of Type I X-ray bursts from the low-mass X-ray binary EXO 0748-676, based on detailed NuSTAR observations, provides critical insights into the mechanisms behind thermonuclear explosions on neutron stars. This study examined the light curve data, revealing the dynamic nature of the accretion process and the conditions necessary for hydrogen and helium ignition. A prominent Type I X-ray burst observed at , with a peak rate of 1,615.88 counts per second, highlights the intense thermonuclear activity caused by the unstable burning of accreted material on the neutron star’s surface.
The findings demonstrate that the thermonuclear ignition is most pronounced in mid-energy bands (7-18 keV), while the cooling phase is dominated by soft X-rays (3-7 keV), underscoring the energy-dependent nature of these bursts. The intervals between bursts, including a 984-second gap between the primary and secondary bursts, provide insights into the accretion dynamics and material instabilities. Additionally, the correlation between the timing of burst events and the neutron star’s optical period suggests a close relationship between accretion from the companion star and burst modulation.
The analysis of the eclipse profile of EXO 0748-676 revealed that higher-energy photons penetrate surrounding material more effectively, as indicated by energy-dependent ingress and egress durations. Variations in and across energy bands reflect the density structure of the environment, revealing complex interactions between X-ray emissions and the material surrounding the neutron star. The study also applied two spectral continuum models to describe the system’s X-ray emissions. Model 1 focused on Comptonization in a high-temperature electron plasma, while Model 2 introduced a dual-component approach, emphasizing both thermal and non-thermal processes. The close agreement between the models strengthens the reliability of the analysis, though differences point to complexities in the emission mechanisms.
When compared to earlier observations, this study suggests temporal changes in the accretion environment of EXO 0748-676. The NuSTAR data revealed a harder spectrum and lower cutoff energies, indicating a cooler Comptonizing region during the observation period. These findings emphasize the dynamic nature of the system and its variable emission characteristics.
In the broader context of neutron star behavior, the results support the idea that neutron stars in low-mass X-ray binaries, such as EXO 0748-676, could exhibit characteristics similar to millisecond pulsars. These neutron stars often rotate rapidly and may emit periodic pulses of radiation due to interactions between their magnetic fields and accretion flows. Although the primary focus of this study was on X-ray bursts and spectral characteristics, the presence of strong magnetic fields and the potential for pulsar-like emissions are important factors in understanding the behavior of such systems.
Overall, this comprehensive study deepens the understanding of X-ray bursts, accretion dynamics, and spectral properties in neutron star systems. Future research, including more pulsar-focused studies, will be essential for further exploring the interactions between accretion, nuclear burning, and the neutron star’s magnetic environment. These investigations will contribute to the broader understanding of neutron star evolution, stellar processes, and high-energy astrophysical phenomena.